# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ C.1 The conformal p-values
$\mathbb{P}\big{(}p_{j}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,\text{ and }\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,$
for any $t\in[0,1]$.
Furthermore, there is $p_{j}\leq\bar{p}_{j}$ and $p_{j}^{\textnormal{rand}}\leq\bar{p}_{j}^{\textnormal{rand}}$ deterministically, where the equalities hold for the nonformity score $S(x,y)=\widehat{\mu}(X)-\big{(}\sup_{x}2|\widehat{\mu}(x)|\big{)}\cdot y$.
Proof.
For any $j\in[m]$ and $t\in[0,1]$, it can be checked that when $Y_{n+j}\leq c_{n+j}$,
@@
where $k^{*}=\min(|{\mathcal{N}}_{j}|,\lfloor t(n+1)\rfloor)$. Let $T_{k^{*}}$ denote the value of the $k^{*}$-th largest element in ${\mathcal{N}}_{j}$. Then apparently, both $k^{*}$ and $T_{k^{*}}$ are invariant to the permutation on ${\mathcal{I}}_{j}$. Consequently,
@@
Using these notations,
$\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}=\mathbb{P}\big{(}U_{n+j}E_{n+j}+G_{n+j}^{+}\leq(n+1)t,Y_{n+j}\leq c_{n+j}\big{)}$ (49)
@@
where the second step follows from exchangeability. Let $i^{*}$ denote the index of the largest element in $\{\widehat{S}_{i}:i\in{\mathcal{N}}_{j},G_{i}^{+}<(n+1)t\}$ (with ties broken arbitrarily). Then
@@
($=\leavevmode\nobreak$ $\frac{1}{n+1}\mathbb{E}\Big{[}\mathds{1}\big{\{}|{\mathcal{N}}_{j}|\leq(n+1)t\big{\}}|{\mathcal{N}}_{j}|+\mathds{1}\{|{\mathcal{N}}_{j}|>(n+1)t\}\cdot\Big{(}G_{i^{*}}^{+}+E_{i^{*}}\cdot\mathds{1}\Big{\{}U_{n+j}\leq\frac{(n+1)t-G_{i^{*}}^{+}}{E_{i^{*}}}\Big{\}}\Big{)}\Big{]}$
$=\frac{1}{n+1}\mathbb{E}\Big{[}\mathds{1}\big{\{}|{\mathcal{N}}_{j}|\leq(n+1)t\big{\}}(n+1)t+\mathds{1}\{|{\mathcal{N}}_{j}|>(n+1)t\}\cdot(n+1)t\Big{]}$
$\leq t.$
By definition,
$p_{j}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{Y_{i}\leq c_{i},\widehat{S}_{i}\geq\widehat{S}_{n+j}\}}{n+1}$ $\overset{\rm(a)}{\leq}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{Y_{i}\leq c_{i},S_{i}\geq\widehat{S}_{n+j}\}}{n+1}$
$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}=\bar{p}_{j}.$
@@
Above, the step (a) is because when $Y_{i}\leq c_{i}$, $S_{i}\geq\widehat{S}_{i}$; when $S(x,y)=\widehat{\mu}(x)-(\sup_{x}2|\widehat{\mu}(x)|)\cdot\mathds{1}\{y>c\}$, the $\leq$” is replaced by “$=$” in step (a) and (b). ∎

### v3

@@ E.2 The conformal p-values
$\mathbb{P}\big{(}p_{j}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,\quad\text{and}\quad\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,\quad\forall t\in[0,1].$
Furthermore, there is $p_{j}\leq\bar{p}_{j}$ and $p_{j}^{\textnormal{rand}}\leq\bar{p}_{j}^{\textnormal{rand}}$ deterministically. If the selection score $S(x,y)=\widehat{\mu}(X)-\big{(}\sup_{x}2|\widehat{\mu}(x)|\big{)}\cdot y$, where $\widehat{\mu}$ is regression function fitted on ${\mathcal{D}}_{\textnormal{calib}}$, then the equalities hold.
Proof of Lemma 2.
For any $j\in[m]$ and $t\in(0,1)$, it can be checked that when $Y_{n+j}\leq c_{n+j}$,
@@
where $k^{*}=\min(|{\mathcal{N}}_{j}|,\lfloor t(n+1)\rfloor)$. Let $T_{k^{*}}$ denote the value of the $k^{*}$-th largest element in ${\mathcal{N}}_{j}$. Apparently, both $k^{*}$ and $T_{k^{*}}$ are invariant to the permutation on ${\mathcal{I}}_{j}$. Consequently,
@@
Using these notations, we can write
$\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}=\mathbb{P}\big{(}U_{n+j}E_{n+j}+G_{n+j}^{+}\leq(n+1)t,Y_{n+j}\leq c_{n+j}\big{)}$
@@
where the first step follows from exchangeability. Let $i^{*}$ denote the index of the largest element in $\{\widehat{S}_{i}:i\in{\mathcal{N}}_{j},G_{i}^{+}<(n+1)t\}$ (with ties broken arbitrarily). Then
@@
($=\leavevmode\nobreak$ $\frac{1}{n+1}\mathbb{E}\Bigg{[}\mathds{1}\big{\{}|{\mathcal{N}}_{j}|\leq(n+1)t\big{\}}\cdot|{\mathcal{N}}_{j}|+\mathds{1}\{|{\mathcal{N}}_{j}|>(n+1)t\}\cdot\bigg{(}G_{i^{*}}^{+}+E_{i^{*}}\cdot\mathds{1}\Big{\{}U_{n+j}\leq\frac{(n+1)t-G_{i^{*}}^{+}}{E_{i^{*}}}\Big{\}}\bigg{)}\Bigg{]}$
[…]
We have therefore proved the validity of $p_{j}$’s and $p_{j}^{\textnormal{rand}}$’s.
Next, we have
$p_{j}=\frac{1+\sum_{i\in[n]}\mathds{1}\{Y_{i}\leq c_{i},\widehat{S}_{i}\geq\widehat{S}_{n+j}\}}{n+1}$ $\overset{\rm(a)}{\leq}\frac{1+\sum_{i\in[n]}\mathds{1}\{Y_{i}\leq c_{i},S_{i}\geq\widehat{S}_{n+j}\}}{n+1}$
$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in[n]}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}=\bar{p}_{j}.$
@@
Above, the step (a) is because when $Y_{i}\leq c_{i}$, $S_{i}\geq\widehat{S}_{i}$; when $S(x,y)=\widehat{\mu}(x)-(\sup_{x}2|\widehat{\mu}(x)|)\cdot\mathds{1}\{y>c\}$, the $\leq$” is replaced by “$=$” in step (a) and (b). A similar argument leads to that results corresponding to $p_{j}^{\textnormal{rand}}$. ∎
E.3 Additional results on the conformal p-value-based procedures
Below, Lemma 3 and Lemma 4 show that that fixed-threshold and BH-threshold selection rules are special instances of the general selection rule described in Section 4.2 of the main text.
Lemma 3.
Let $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]:p_{j}\leq q\}$. Then $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$ is equivalent to ${\widetilde{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]:\widehat{{\mathcal{S}}}_{n+j}\geq T\}$, where
$T=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}\geq q(n+1)\bigg{\}}.$
Proof of Lemma 3.
Suppose $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$, then $p_{j}\leq q$. By the definition of $p_{j}$,
$\frac{1}{n+1}\Big{(}1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}=p_{j}\leq q.$
By the definition of $T$, $\widehat{S}_{n+j}\geq T$, and therefore $j\in{\widetilde{\mathcal{S}}}_{\textnormal{fixed}}$. Conversely, if $\widehat{S}_{n+j}\geq T$, then
$p_{j}=\frac{1}{n+1}\Big{(}1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}\leq\Big{(}1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq T,Y_{i}\leq c_{i}\}\Big{)}\leq q,$
where the last step is by the definition of $T$. As a result, $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$. We then have $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}={\widetilde{\mathcal{S}}}_{\textnormal{fixed}}$. ∎
Lemma 4.
Let $\widehat{{\mathcal{S}}}_{\textnormal{BH}}$ denote the selection set obtained by the BH procedure applied to the conformal p-values. Then $\widehat{{\mathcal{S}}}_{\textnormal{BH}}$ is equivalent to ${\widetilde{\mathcal{S}}}_{\textnormal{BH}}:=\{j\in[m]:\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$, where
$T^{\textnormal{BH}}=\inf\bigg{\{}t\in{\mathcal{T}}:\frac{1+\sum_{i\in[n]}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum_{j\in[m]}\mathds{1}\{\widehat{S}_{n+j}\geq t\})\vee 1}\frac{m}{1+n}\leq q\bigg{\}}.$
Proof of Lemma 4.
Let $j^{*}=\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$ and $\widehat{S}_{n+(1)},\widehat{S}_{n+(2)},\ldots,\widehat{S}_{n+(m)}$ denote the order statistics of $\{\widehat{S}_{n+j}\}_{j\in[m]}$ in a descending order. It then suffices to show that $j^{*}=k^{*}$. When $k^{*}\geq 1$,
$\frac{1+\sum_{i\in[n]}\{\widehat{S}_{i}\geq\widehat{S}_{n+(k^{*})},Y_{i}\leq c_{i}\}}{(\sum_{j\in[m]}\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{n+(k^{*})}\})\vee 1}\frac{m}{1+n}=p_{(k^{*})}\frac{m}{k^{*}}\leq q,$ (50)
$p_{(j^{*})}\frac{m}{j^{*}}$ =$\overset{\rm(a)}{=}\frac{1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+(j^{*})},Y_{i}\leq c_{i}\}}{n+1}\cdot\frac{m}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}$
$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{BH}},Y_{i}\leq c_{i}\}}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}\cdot\frac{m}{n+1}\overset{\rm(c)}{\leq}q.$
$\frac{1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq t\})\vee 1}\frac{m}{n+1}=\frac{1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{j(t)}\frac{m}{n+1}$
[…]
where the last step is because $k^{*}=0$. The above implies that $T^{\textnormal{BH}}>\widehat{S}_{n+(1)}\Rightarrow j^{*}=0$. The equivalence is therefore established. ∎
E.4 Prediction set construction by further sample splitting
In this section, we discuss the possibility of simplifying the prediction set construction by further splitting the calibration set. Since we are primarily interested in the case where we do not have the flexibiility to change the selection rule, it turns out that—in most of the cases—the sample-splitting idea does not help with finding the exchangeable reference set without substantially changing the nature of selection.
There is one class of selection rules, however, that can benefit from the sample-splitting idea with only minor modifications to the original selection rule. To be concrete, consider the selection rule such that $\mathds{1}\{j\in\widehat{\mathcal{S}}\}=f({\mathcal{D}}_{\textnormal{calib}},X_{n+j})$ for some function $f$, i.e., whether $j$ is selected depends solely on the calibration set ${\mathcal{D}}_{\textnormal{calib}}$ and the feature $X_{n+j}$. Examples of such selection rules include thresholding conformal p-values with a fixed threshold, and selecting units based on properties of the preliminary prediction set.
For such selection rules, we could split the calibration set into two parts, ${\mathcal{D}}_{\textnormal{calib}}={\mathcal{D}}_{\textnormal{calib}}^{(1)}\cup{\mathcal{D}}_{\textnormal{calib}}^{(2)}$. If we allow ourselves to determine the selection set via $\mathds{1}\{j\in\widehat{\mathcal{S}}\}=f({\mathcal{D}}_{\textnormal{calib}}^{(1)},X_{n+j})$, then we can construct the prediction set as
$\begin{split}&\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}:V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}}\cup\{\infty\},1-\alpha\big{)}\Big{\}},\\ &\text{where }\widehat{{\mathcal{R}}}:=\{Z_{i}\in{\mathcal{D}}_{\textnormal{calib}}^{(2)}:f({\mathcal{D}}_{\textnormal{calib}}^{(1)},X_{i})=1\}.\end{split}$ (51)
The following proposition shows that the prediction set constructed in (51) is a valid $(1-\alpha)$ selection-conditional prediction set.
Proposition 8.
Under the condition of Theorem 1, and suppose that for $\forall j\in[m]$, the selection rule is such that $\mathds{1}\{\ell\in\widehat{\mathcal{S}}\}=f({\mathcal{D}}_{\textnormal{calib}},X_{n+\ell})$ for some function $f$. Then the prediction set $\widehat{C}_{\alpha,n+j}$ constructed in (51) is a valid $(1-\alpha)$ prediction set.
Proof.
Treating ${\mathcal{D}}_{\textnormal{calib}}^{(2)}$ as the calibration set in Theorem 1, it is straight forward to check that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}$, and therefore by Theorem 1, the prediction set $\widehat{C}_{\alpha,n+j}$ is a valid $(1-\alpha)$ prediction set. ∎
E.5 Analysis of calibration sample size
A practical concern of our proposed method is that accounting for the selection event may reduce the number of calibration samples, which has a direct impact on the size and stability of the prediction set. The degree of effect is high dependent on the selection rule, which determines the fundamental difficulty of the problem: for example, if the selection rule picks out only the test unit with the largest predicted value, then conditional inference essentially becomes inference on the tail of a distribution; it is not surprising to see that only a limited number of calibration samples can be obtained.
In this section, we provide an asymptotic analysis of a class of selection rules that are relatively “mild”. To be concrete, we consider the selection rule yields a reference set in the form of $\{i\in[n]:h(X_{i})\geq\widehat{q}\}$, where $h$ is a function of the features and $\widehat{q}$ is a threshold that potentially dependent on $({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ but converges to a constant as $n,m$ goes to infinity. The following proposition characterizes the size of such reference set asymptotically.
Proposition 9.
Suppose the reference set ^=$\widehat{{\mathcal{R}}}=\{i\in[n]:h(X_{i})\geq\widehat{q}\}$, where $h$ is a fixed function of the features and $\widehat{q}$ is a data-dependent threshold such that $\lim_{n,m\rightarrow}\widehat{q}=q^{*}$ a.s. For $\forall t\in\mathbb{R}$, let $G(t)=\mathbb{P}(h(X)\geq t)$. Assume that $G(t)$ is continuous in $t$. Then the size of the reference set satisfies that
[…]
Proof.
By definition, we can write |^$|\widehat{{\mathcal{R}}}|=\sum_{i=1}^{n}\mathds{1}\{h(X_{i})\geq\widehat{q}\}$. By the strong law of large numbers, we have that
[…]
Since $\lim_{n,m\rightarrow\infty}\widehat{q}=q^{*}$, and $G(t)$ is continuous in $t$, we have by the continuous mapping theorem that $\lim_{n,m\rightarrow\infty}|G(\widehat{q})-G(q^{*})|=0$. By triangular inequality, we complete the proof. ∎
Many selection rules we have considered in this paper fall into the class. For example, for the quantile based selection rule, the reference set is $\{i\in[n]:S(X_{i})\geq\widehat{q}\}$, where $\widehat{q}$ is either the joint quantile or the calibration quantile, which converges to the popuplation quantile as $n,m$ goes to infinity. The class of conformal p-value-based selection rules considered in Section 4.2 can also be characterized by the proposition. To see this, consider the fixed threshold selection rule or the BH-threshold selection rule. For simplicity suppose that $y\leq c_{n+j}$ and $\mathfrak{S}=2^{[m]}$, then
^$\widehat{{\mathcal{R}}}_{n+j}(y)=\big{\{}i\in[n]:Y_{i}\leq c_{i},\widehat{S}_{i}\geq\tau^{(j)}(1,0)\big{\}}\cup\big{\{}i\in[n]:Y_{i}>c_{i},\widehat{S}_{i}\geq\tau^{(j)}(1,1)\big{\}}.$
Above, $\tau^{(j)}(1,0)$ and $\tau^{(j)}(1,1)$ converge to the same constant for both variants (c.f. Jin and Candès, 2023b ).